1. As you start consuming potato chips, your marginal utility is very high, but it begins to fall slowly until you've eaten 10 chips. After you have eaten 10 chips, marginal utility decreases even faster with each additional chip. Marginal utility continues to decline until you’ve eaten 49 chips. The fiftieth chip does not give you any additional utility. After 50 chips, your mouth gets so salty that it is unpleasant to eat any more, so marginal utility is actually negative for those chips. How many chips should you eat in order to maximize your marginal utility?

1

10

49

50

2. At current levels of consumption, Alice is spending her entire budget. If Alice gets 3 utils of satisfaction per dollar spent on ice cream and 2 utils per dollar spent on shampoo, then how should she adjust her consumption to get closer to the consumer optimum?

She is already at the consumer optimum, so no adjustment is necessary.

She should buy more ice cream and less shampoo.

She should buy more shampoo and less ice cream.

She should buy more of both goods.

3. Suppose you are at a restaurant and your favorite dish costs $20. You are willing to pay up to $17 for your next favorite dish. If your next-favorite dish gives you 100 utils, how many additional utils do you need from your favorite dish to spend the extra $3?

15

17.65

85

117.65

4. Suppose you plan on eating 50 potato chips. As you start consuming potato chips, your marginal utility is very high, but it begins to fall slowly until you’ve eaten 10 chips. After you have eaten 10 chips, your marginal utility decreases even faster with each additional chip. Marginal utility continues to decline until you’ve eaten 49 chips. The fiftieth chip does not give you any additional utility. After 50 chips, your mouth gets so salty that it is unpleasant to eat any more, so marginal utility is actually negative for those chips. How many chips should you eat in order to maximize your total utility?

1

10

49

51
Answer #1

Ans.1- 1

1st unit will give maximum marginal utility due to law of diminishing marginal returns.

Ans.2- She should buy more ice cream and less shampoo.

She should buy more icecream because marginal utility per dollar spent is higher for ice cream.

Ans.3- 17.65

Marginal utility per dollar spent = 100/17 = 5.88

Additional utils from $3 = 3*5.88 = 17.65

Ans.4- 49